# Lusztig sheaves, characteristic cycles and the Borel-Moore homology of Nakajima’s quiver varieties

### v1

@@ 3.2. Induction and restriction functors
3.3. Fourier-Laumon transforms
For $h\in\Omega$, let $(I,H,\Omega_{h})$ be the quiver obtained from the quiver $Q=(I,H,\Omega)$ by reversing the orientation of $h$, that is, $\Omega_{h}=(\Omega\backslash\{h\})\cup\{\bar{h}\}$. Following [1, Section 10.4], there is the Fourier-Laumon transform functor
${\rm{Four}}_{h}:\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega_{h}})$
which is a derived equivalence and commutes with the induction functor
$\mathbf{Ind}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}({\rm{Four}}_{h}(L)\boxtimes{\rm{Four}}_{h}(L^{\prime}))={\rm{Four}}_{h}(\mathbf{Ind}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}(L\boxtimes L^{\prime})),$
see [1, Proposition 10.4.5]. In general, for any other orientation $\Omega^{\prime}\subset H$, we can define the Fourier-Laumon transform functor
${\rm{Four}}_{\Omega,\Omega^{\prime}}:\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega^{\prime}})$
by composing those ${\rm{Four}}_{h},h\in\Omega\cap\Omega^{\prime}$ in a fixed order.
@@ 3.3. Fourier-Laumon transforms
We remark that Lusztig used the Fourier-Deligne transform in [14] and [16], because he considered $\overline{\mathbb{F}}_{q}$-varieties and ¯$\overline{\mathbb{Q}}_{l}$-sheaves. In this paper, we consider $\mathbb{C}$-varieties and $\mathbb{C}$-sheaves, and so the Fourier-Deligne transform should be replaced by the Fourier-Laumon transform. The statements in [14] and [16] involving the Fourier-Deligne transform still hold after replacing by the Fourier-Laumon transform.
@@ 3.4.1. Analysis at sink
Let $i$ be a sink for the orientation $\Omega$, that is, there is no $h\in\Omega$ such that $h^{\prime}=i$. For any $\nu\in\mathbb{N}I,0\leqslant t\leqslant\nu_{i}$, we define ${{}^{t}\mathbf{E}_{\mathbf{V},\Omega}^{i}}\subset\mathbf{E}_{\mathbf{V},\Omega}$ to be the locally closed subset consisting of $x$ such that
@@
then all these ${{}^{t}\mathbf{E}_{\mathbf{V},\Omega}^{i}},0\leqslant t\leqslant\nu_{i}$ form a $(\mathbf{G}_{\mathbf{V}}\times\mathbb{T})$-invariant stratification of $\mathbf{E}_{\mathbf{V},\Omega}$. For any simple perverse sheaf $L\in\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega})$, there exists a unique integer $t_{i}(L)$ such that ${{}^{t_{i}(L)}\mathbf{E}_{\mathbf{V},\Omega}^{i}}\cap\mathrm{supp}(L)$ is open dense in $\mathrm{supp}(L)$.
Lemma 3.3 ([14, Lemma 6.4]).
@@ 3.4.2. Analysis at source
Let $i$ be a source for the orientation $\Omega$, that is, there is no $h\in\Omega$ such that $h^{\prime\prime}=i$. For any $\nu\in\mathbb{N}I,0\leqslant t\leqslant\nu_{i}$, we define $\mathbf{E}_{\mathbf{V},\Omega}^{i,t}\subset\mathbf{E}_{\mathbf{V},\Omega}$ to be the locally closed subset consisting of $x$ such that
@@
then all these $\mathbf{E}_{\mathbf{V},\Omega}^{i,t},0\leqslant t\leqslant\nu_{i}$ form a $(\mathbf{G}_{\mathbf{V}}\times\mathbb{T})$-invariant stratification of $\mathbf{E}_{\mathbf{V},\Omega}$. For any simple perverse sheaf $L\in\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega})$, there exists a unique integer $t_{i}^{\ast}(L)$ such that $\mathbf{E}_{\mathbf{V},\Omega}^{i,t_{i}^{\ast}(L)}\cap\mathrm{supp}(L)$ is open dense in $\mathrm{supp}(L)$. The following lemma is dual to Lemma 3.3.
@@ 3.5. Categorification of ${{}_{\mathcal{A}}\mathbf{U}^{-}_{v}}$ and canonical bases
sending [¯$[\overline{\mathbb{Q}}_{l}|_{\mathbf{E}_{\mathbf{V},\Omega}}]$ to $F_{i}^{(n)}$, where $|\mathbf{V}|=ni$ for $i\in I,n\in\mathbb{N}$.
@@ 3.6. Refine string order
For any $\nu\in\mathbb{N}I,L\in\mathcal{P}_{\nu}$ and $i\in I$, we define $0\leqslant\epsilon_{i}(L)\leqslant\nu_{i}$ to be the maximum integer $t$ such that $L$ is a direct summand of $L_{\boldsymbol{\nu}}$ up to shift, where $\boldsymbol{\nu}=(ti,...)\in\mathcal{V}_{\nu}$ is a sequence starting by $ti$. Dually, we define $0\leqslant\epsilon^{\ast}_{i}(L)\leqslant\nu_{i}$ to be the maximum integer $t$ such that $L$ is a direct summand of $L_{\boldsymbol{\nu}}$ up to shift, where $\boldsymbol{\nu}=(...,ti)\in\mathcal{V}_{\nu}$ is a sequence ending by $ti$. If $i$ is a sink for the orientation $\Omega$, we have $t_{i}(L)=\epsilon_{i}(L)$ by [14, Proposition 6.6]. Dually, if $i$ is a source for the orientation $\Omega$, we have $t^{\ast}_{i}(L)=\epsilon^{\ast}_{i}(L)$. If $\nu\not=0$, there exists $i\in I$ such that $\epsilon_{i}(L)>0$ by [14, Lemma 7.1]. Dually, if $\nu\not=0$, there exists $i\in I$ such that $\epsilon^{\ast}_{i}(L)>0$. Then we define two maps $\epsilon_{i},\epsilon^{\ast}_{i}:\mathcal{B}\rightarrow\mathbb{N}$ by
@@
$\bullet$ if $\nu\not=0$, let $i\in I$ be the minimal vertex such that $\epsilon_{i}(L)>0$, then $s^{\prec}(\chi\mathrm{Four}_{\Omega_{i},\Omega}\pi_{i,\epsilon_{i}(L)}\mathrm{Four}_{\Omega,\Omega_{i}}(L))$ has been defined by inductive hypothesis, where $\Omega_{i}$ is an orientation such that $i$ is a sink for it, $\mathrm{Four}_{\Omega_{i},\Omega},\mathrm{Four}_{\Omega,\Omega_{i}}$ are the Fourier-Laumon (or Fourier-Deligne) transforms and $\pi_{i,\epsilon_{i}(L)}$ is the bijection in Lemma 3.3, and we define
@@
We remark that above partial order on $\mathcal{B}$ is defined in [6] and is a refinement of the string order defined in [3, Section 2.5].
@@ 4.1. Lusztig sheaves on the framed quivers and the localizations
which have $\mathbf{G}_{\mathbf{V}\oplus\mathbf{W}^{1}}\times\mathbb{T}$-action and $\mathbf{G}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}\times\mathbb{T}$-action respectively. Throughout this paper, we only consider the $\mathbf{G}_{\mathbf{V}}\times\mathbb{T}$-actions on them. There are corresponding categories $\mathcal{Q}_{\nu+\omega^{1}},\mathcal{Q}_{\nu+\omega^{1}+\omega^{2}}$ of Lusztig sheaves on them respectively, see subsection 3.1.
Suppose $I=\{i_{1},i_{2},\cdots,i_{n}\}$, we fix two sequences
@@
Proposition 4.2 ([7, Lemma 4.3, Corollary 4.7]).
@@
We only need to replace the Fourier-Deligne transform by the Fourier-Laumon transform in the proof of [7, Lemma 4.3, Corollary 4.7]. ∎
@@
(b) For any $i\in I$, we fix an orientation $\Omega_{i}$ such that $i$ is a source, then there are thick subcategories $\mathcal{N}_{\nu,i}$ of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega_{i}^{(1)}})$ and of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega_{i}^{(2)}})$ by (a). We define $\mathcal{N}_{\nu}$ to be the thick subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}})$ generated by ${\rm{Four}}_{\Omega_{i}^{(1)},\Omega^{(1)}}(\mathcal{N}_{\nu,i})$ for any $i\in I$, where ${\rm{Four}}_{\Omega_{i}^{(1)},\Omega^{(1)}}$ is Fourier-Laumon transform for the framed quiver $Q^{(1)}$. Similarly, we define $\mathcal{N}_{\nu}$ to be the thick subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})$ generated by ${\rm{Four}}_{\Omega_{i}^{(2)},\Omega^{(2)}}(\mathcal{N}_{\nu,i})$ for any $i\in I$, where ${\rm{Four}}_{\Omega_{i}^{(2)},\Omega^{(2)}}$ is Fourier-Laumon transform for the two-framed quiver $Q^{(2)}$.

### v3

@@ 3.2. Induction and restriction functors
3.3. Fourier-Sato transforms
In this section, we recall the definition of Fourier-Sato transform for quivers in [13]. One can also see details in [2] and [17] for general manifolds. Assume $i$ is a source in $\Omega$ , that is, there is no $h\in\Omega$ such that $h^{\prime\prime}=i$, then there is a natural projection
$\pi_{i}:\mathbf{E}_{\mathbf{V},\Omega}\rightarrow\dot{\mathbf{E}}_{\mathbf{V},i}=\bigoplus_{h\in\Omega,h^{\prime}\neq i}{\rm{Hom}}(\mathbf{V}_{h^{\prime}},\mathbf{V}_{h^{\prime\prime}}).$
The formula $\lambda:\mathbb{G}_{m}\rightarrow\mathbf{G}_{\mathbf{V}},t\mapsto t^{-1}\textrm{Id}_{\mathbf{V}_{i}}$ defines an action of $\mathbb{G}_{m}$ on the fiber of $\pi_{i}$, which contracts the fiber. Similarly, let $\Omega^{\prime}=\mu_{i}(\Omega)$ be the orientation obtained from $\Omega$ by reversing all arrows starting from $i$, then $i$ is a sink in $\Omega^{\prime}$, that is, there is no $h\in\Omega^{\prime}$ such that $h^{\prime}=i$. We can consider the action of $\lambda^{\prime}:\mathbb{G}_{m}\rightarrow\mathbf{G}_{\mathbf{V}},t\mapsto t\textrm{Id}_{\mathbf{V}_{i}}$ on $\mathbf{E}_{\mathbf{V},\Omega^{\prime}}$, then $\mathbf{E}_{\mathbf{V},\Omega^{\prime}}$ is identified with the dual bundle of $\mathbf{E}_{\mathbf{V},\Omega}$ over $\dot{\mathbf{E}}_{\mathbf{V},i}$ via the trace form.
${\mathbf{E}_{\mathbf{V},\Omega}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{i}}$${\mathbf{E}_{\mathbf{V},\Omega^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{{}^{\prime}}_{i}}$${\dot{\mathbf{E}}_{\mathbf{V},i}}$
Notice that any object in $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$ is monodromic with respect to $\lambda$ in the sense of [2], by [13, D.4] the Fourier-Sato transform defines a perverse equivalence
${\rm{Four}}_{i}:\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega^{\prime}}).$
Since Fourier-Sato transforms are compatible with pull-back or push-forward of base change and morphisms between vector bundles, one can apply the proof of [1, 10.4.5] for Fourier-Laumon transforms, and prove that ${\rm{Four}}_{i}$ commutes with the induction functor
$\mathbf{Ind}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}({\rm{Four}}_{i}(L)\boxtimes{\rm{Four}}_{i}(L^{\prime}))={\rm{Four}}_{i}(\mathbf{Ind}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}(L\boxtimes L^{\prime})).$
In general, for any other orientation $\Omega,\Omega^{\prime}\subset H$ without oriented cycles, we assume that there is a sequence $(i_{1},i_{2},\cdots,i_{m})$ of $i\in I$, such that $i_{s}$ is a sink in $\mu_{s-1}\cdots\mu_{1}(\Omega)$ for any $s$, and $\mu_{m}\cdots\mu_{1}(\Omega)=\Omega^{\prime}$. Consider the vector bundle $\mathbf{E}_{\mathbf{V},\Omega}\rightarrow\mathbf{E}_{\mathbf{V},\Omega\cap\Omega^{\prime}}$ and define the corresponding Fourier-Sato transform ${\rm{Four}}_{\Omega,\Omega^{\prime}}$. In this case, we say $\Omega,\Omega^{\prime}$ are equivalent under mutations.
Lemma 3.3.
With the above assumption, any object in $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$ is monodromic with respect to the above vector bundle and the Fourier-Sato transform is well-defined on $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$.
Proof.
We prove that there exists an one parameter subgroup $\lambda:\mathbb{G}_{m}\rightarrow\mathbf{G}_{\mathbf{V}}$ such that $\lambda(\mathbb{G}_{m})$ acts on $\mathbf{E}_{\mathbf{V},\Omega}$ by contracting the fiber of $\mathbf{E}_{\mathbf{V},\Omega}\rightarrow\mathbf{E}_{\mathbf{V},\Omega\cap\Omega^{\prime}}$ by induction on the length $m$ of the sequence. When $m=1$, we use the action of $\mathbb{G}_{m}$ for ${\rm{Four}}_{i_{1}}$. If $\lambda_{k}:\mathbb{G}_{m}\rightarrow\mathbf{G}_{\mathbf{V}}$ has been constructed for $(i_{1},i_{2},\cdots,i_{k})$, then we can take $\lambda_{k+1}(t)=\lambda_{k}(t)\circ t\textrm{Id}_{\mathbf{V}_{k+1}}$.
Let $\Omega_{k}=\mu_{k}\cdots\mu_{1}(\Omega)$, then $\lambda_{k}(t)$ acts on $x_{h}$ of arrows $h\in\Omega\backslash(\Omega\cap\Omega_{k})$ with weight $1$. Notice that if $h\in\Omega\backslash(\Omega\cap\Omega_{k})$ and $h$ is adjacent to $k+1$, then $\bar{h}\in\Omega_{k}$ must has a target $k+1$, since $k+1$ is a sink in $\Omega_{k}$. In particular, $h$ starts at $k+1$ and $t\textrm{Id}_{\mathbf{V}_{k+1}}$ acts on such $x_{h}$ by $t^{-1}$. Hence $\lambda_{k+1}(t)$ acts trivially on such $x_{h}$. When $h\in\Omega\backslash(\Omega\cap\Omega_{k})$ is not adjacent to $k+1$, $\lambda_{k+1}(t)$ acts on such $x_{h}$ as $\lambda_{k}(t)$ with weight $1$. If $h\notin\Omega\backslash(\Omega\cap\Omega_{k})$, then $\lambda_{k+1}(t)$ acts on such $x_{h}$ as $t\textrm{Id}_{\mathbf{V}_{k+1}}$. With the argument above, we can see that $\lambda_{k+1}(t)$ acts by contracting $x_{h}$ with $h\in\Omega\backslash(\Omega\cap\Omega_{k+1}).$
With the action of $\lambda:\mathbb{G}_{m}\rightarrow\mathbf{G}_{\mathbf{V}}$, any object in $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$ is $\mathbb{G}_{m}$-equivariant, hence is monodromic. ∎
Lemma 3.4.
For any acyclic orientation $\Omega$ and $i\in I$, there exists an orientation $\Omega_{i}$ equivalent to $\Omega$ under mutations such that $i$ is a source in $\Omega_{i}$. Similarly, there exists an orientation $\Omega^{i}$ equivalent to $\Omega$ under mutations such that $i$ is a sink in $\Omega^{i}$.
Proof.
We only prove the first statement. Consider the set $P(\Omega,i)$ of paths in $Q$ with target $i$. Since there is no oriented cycle in $\Omega$, there exists a minimal integer $m$ such that all paths in $P(\Omega,i)$ has length $\leqslant m$. If $m=0$, then $i$ is a source in $\Omega$. If $m>0$, then we consider the paths in $P(\Omega,i)$ with length equal to $m$ and denote their sources by $i_{1},i_{2},\cdots,i_{t}\in I$. Then $i_{1},i_{2},\cdots,i_{t}$ must be sources in $\Omega$, otherwise, there exists a path with length $>m$ in $\Omega$. After applying $\mu_{i_{1}},\cdots,\mu_{i_{t}}$, we obtain an orientation $\tilde{\Omega}$ such that paths in $P(\tilde{\Omega},i)$ has length $\leqslant m-1$. By induction on $m$, we get the sequence desired. ∎
@@ 3.3. Fourier-Sato transforms
We remark that Lusztig used the Fourier-Deligne transform in [18] and [20], because he considered $\overline{\mathbb{F}}_{q}$-varieties and $\mathbb{C}$-sheaves. In this paper, we consider $\mathbb{C}$-varieties and $\mathbb{C}$-sheaves, and so the Fourier-Deligne transform should be replaced by the Fourier-Sato transform. The statements in [18] and [20] involving the Fourier-Deligne transform still hold after replacing by the Fourier-Sato transform.
@@ 3.4. Lusztig’s key lemmas
By Lemma 3.3 and 3.4, we can always apply the Fourier-Sato transform and assume $i$ is a source or sink.
@@ 3.4.1. Analysis at sink
Let $i$ be a sink for the orientation $\Omega$. For any $\nu\in\mathbb{N}I,0\leqslant t\leqslant\nu_{i}$, we define ${{}^{t}\mathbf{E}_{\mathbf{V},\Omega}^{i}}\subset\mathbf{E}_{\mathbf{V},\Omega}$ to be the locally closed subset consisting of $x$ such that
@@
then all these ${{}^{t}\mathbf{E}_{\mathbf{V},\Omega}^{i}},0\leqslant t\leqslant\nu_{i}$ form a $\mathbf{G}_{\mathbf{V}}$-invariant stratification of $\mathbf{E}_{\mathbf{V},\Omega}$. For any simple perverse sheaf $L\in\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$, there exists a unique integer $t_{i}(L)$ such that ${{}^{t_{i}(L)}\mathbf{E}_{\mathbf{V},\Omega}^{i}}\cap\mathrm{supp}(L)$ is open dense in $\mathrm{supp}(L)$.
Lemma 3.5 ([18, Lemma 6.4]).
@@ 3.4.2. Analysis at source
Let $i$ be a source for the orientation $\Omega$. For any $\nu\in\mathbb{N}I,0\leqslant t\leqslant\nu_{i}$, we define $\mathbf{E}_{\mathbf{V},\Omega}^{i,t}\subset\mathbf{E}_{\mathbf{V},\Omega}$ to be the locally closed subset consisting of $x$ such that
@@
then all these $\mathbf{E}_{\mathbf{V},\Omega}^{i,t},0\leqslant t\leqslant\nu_{i}$ form a $\mathbf{G}_{\mathbf{V}}$-invariant stratification of $\mathbf{E}_{\mathbf{V},\Omega}$. For any simple perverse sheaf $L\in\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$, there exists a unique integer $t_{i}^{\ast}(L)$ such that $\mathbf{E}_{\mathbf{V},\Omega}^{i,t_{i}^{\ast}(L)}\cap\mathrm{supp}(L)$ is open dense in $\mathrm{supp}(L)$. The following lemma is dual to Lemma 3.5.
@@ 3.5. Categorification of ${{}_{\mathcal{A}}\mathbf{U}^{-}_{v}}$ and canonical bases
sending $[\mathbb{C}|_{\mathbf{E}_{\mathbf{V},\Omega}}]$ to $F_{i}^{(n)}$, where $|\mathbf{V}|=ni$ for $i\in I,n\in\mathbb{N}$.
@@ 3.6. Refine string order
For any $\nu\in\mathbb{N}I,L\in\mathcal{P}_{\nu}$ and $i\in I$, we define $0\leqslant\epsilon_{i}(L)\leqslant\nu_{i}$ to be the maximum integer $t$ such that $L$ is a direct summand of $L_{\boldsymbol{\nu}}$ up to shift, where $\boldsymbol{\nu}=(ti,...)\in\mathcal{V}_{\nu}$ is a sequence starting by $ti$. Dually, we define $0\leqslant\epsilon^{\ast}_{i}(L)\leqslant\nu_{i}$ to be the maximum integer $t$ such that $L$ is a direct summand of $L_{\boldsymbol{\nu}}$ up to shift, where $\boldsymbol{\nu}=(...,ti)\in\mathcal{V}_{\nu}$ is a sequence ending by $ti$. If $i$ is a sink for the orientation $\Omega$, we have $t_{i}(L)=\epsilon_{i}(L)$ by [18, Proposition 6.6]. Dually, if $i$ is a source for the orientation $\Omega$, we have $t^{\ast}_{i}(L)=\epsilon^{\ast}_{i}(L)$. If $\nu\not=0$, there exists $i\in I$ such that $\epsilon_{i}(L)>0$ by [18, Lemma 7.1]. Dually, if $\nu\not=0$, there exists $i\in I$ such that $\epsilon^{\ast}_{i}(L)>0$. Then we define two maps $\epsilon_{i},\epsilon^{\ast}_{i}:\mathcal{B}\rightarrow\mathbb{N}$ by
@@
$\bullet$ if $\nu\not=0$, let $i\in I$ be the minimal vertex such that $\epsilon_{i}(L)>0$, then $s^{\prec}(\chi\mathrm{Four}_{\Omega_{i},\Omega}\pi_{i,\epsilon_{i}(L)}\mathrm{Four}_{\Omega,\Omega_{i}}(L))$ has been defined by inductive hypothesis, where $\Omega_{i}$ is an orientation such that $i$ is a sink for it, $\mathrm{Four}_{\Omega_{i},\Omega},\mathrm{Four}_{\Omega,\Omega_{i}}$ are the Fourier-Sato (or Fourier-Deligne) transforms and $\pi_{i,\epsilon_{i}(L)}$ is the bijection in Lemma 3.5, and we define
@@
We remark that above partial order on $\mathcal{B}$ is defined in [11] and is a refinement of the string order defined in [4, Section 2.5].
@@ 4.1. Lusztig sheaves on the framed quivers and the localizations
which have $\mathbf{G}_{\mathbf{V}\oplus\mathbf{W}^{1}}$-action and $\mathbf{G}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$-action respectively.
We also define the torus
$\mathbb{T}=\mathbb{G}_{m}\times...\times\mathbb{G}_{m}$
of rank $\sharp I$ such that it acts on $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}$ and $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}$ respectively by
$t=(t_{i})_{i\in I}\mapsto(t_{i}\textrm{Id}_{\mathbf{W}^{1}})_{i\in I}\in\mathbf{G}_{\mathbf{V}\oplus\mathbf{W}^{1}},$
$t=(t_{i})_{i\in I}\mapsto(t_{i}\textrm{Id}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}})_{i\in I}\in\mathbf{G}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}.$
Throughout this paper, we only consider the $\mathbf{G}_{\mathbf{V}}\times\mathbb{T}$-actions on $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}$ and $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}$, and the $\mathbf{G}_{\mathbf{V}}\times\mathbb{T}$ equivariant derived categories of constructible sheaves on $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}$ and $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}$. Here we need the $\mathbb{T}$-action, because we need to use the Fourier-Sato transforms in subsection 3.3 to reverse arrows of the form $i\rightarrow i^{1},i\rightarrow i^{2}$ or $i\leftarrow i^{1},i\leftarrow i^{2}$. We denote these mutations reversing framed arrows respectively by $\mu_{i^{1}}$ and $\mu_{i^{1}}\mu_{i^{2}}$ for framed and 2-framed quivers. If two orientations $\Omega^{(1)}$ and $\Omega^{(1),^{\prime}}$ can be obtained by a sequence of $\mu_{i}$ or $\mu_{i^{1}}$, we say they are equivalent under mutation. If two orientations $\Omega^{(2)}$ and $\Omega^{(2),^{\prime}}$ can be obtained by a sequence of $\mu_{i}$ and $\mu_{i^{1}}\mu_{i^{2}}$, we say they are equivalent under mutation. By Lemma 3.3, we can define Fourier-Sato transform for any equivalent orientations.
There are corresponding categories $\mathcal{Q}_{\nu+\omega^{1}},\mathcal{Q}_{\nu+\omega^{1}+\omega^{2}}$ of Lusztig sheaves on them respectively, see subsection 3.1. Suppose $I=\{i_{1},i_{2},\cdots,i_{n}\}$, we fix two sequences
@@
Proposition 4.2 ([10, Lemma 4.3, Corollary 4.7]).
@@
We only need to replace the Fourier-Deligne transform by the Fourier-Sato transform constructed by the above lemma in the proof of [10, Lemma 4.3, Corollary 4.7]. ∎
@@
(b) For any $i\in I$, we fix an orientation $\Omega_{i}$ such that $i$ is a source, then there are thick subcategories $\mathcal{N}_{\nu,i}$ of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega_{i}^{(1)}})$ and of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega_{i}^{(2)}})$ by (a). We define $\mathcal{N}_{\nu}$ to be the thick subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}})$ generated by ${\rm{Four}}_{\Omega_{i}^{(1)},\Omega^{(1)}}(\mathcal{N}_{\nu,i})$ for any $i\in I$, where ${\rm{Four}}_{\Omega_{i}^{(1)},\Omega^{(1)}}$ is Fourier-Sato transform for the framed quiver $Q^{(1)}$. Similarly, we define $\mathcal{N}_{\nu}$ to be the thick subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})$ generated by ${\rm{Four}}_{\Omega_{i}^{(2)},\Omega^{(2)}}(\mathcal{N}_{\nu,i})$ for any $i\in I$, where ${\rm{Four}}_{\Omega_{i}^{(2)},\Omega^{(2)}}$ is Fourier-Sato transform for the two-framed quiver $Q^{(2)}$.